OpenStudy (anonymous):

√((125²)^-⅓) solve without using calculator

OpenStudy (mrnood):

note that 125 = 5^3 also note that sqrt(x) is the same as x^1/2 now use properties of exponents to simplify

OpenStudy (unklerhaukus):

\[\sqrt{(125^2)^{-1/3}}=((125^2)^{-1/3})^{1/2}=125^{(2\times-\tfrac13\times\tfrac12)}\]

OpenStudy (anonymous):

To be polite both solutions are wrong

OpenStudy (anonymous):

i promise to follow the person that can get this answer right

OpenStudy (unklerhaukus):

\[\sqrt{(125^2)^{-1/3}}=((125^2)^{-1/3})^{1/2}=125^{(2\times-\tfrac13\times\tfrac12)}=125^{-1/3}=(5^3)^{-1/3}=...\]

OpenStudy (anonymous):

what is the answer

OpenStudy (unklerhaukus):

\[(5^3)^{-1/3}=5^{3\times-1/3}=...\]

OpenStudy (anonymous):

its wrong

OpenStudy (anonymous):

who needs a clue

OpenStudy (mrnood):

@daniel56K I don't need a clue - I have worked out ht eanswer correctly I have given oyu the correct way to do it hanve YOU worked it out?

OpenStudy (mrnood):

125^2 = 5^3^2 = 5^6 5^6 ^-1/3 = 5 ^-2 5^-2 ^ 1/2 = 5^-1 = 1/5 = 0.2

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

That is correct

OpenStudy (mrnood):

@Daniel56k if oyu look at @UnkleRhaukus answer you will see that he already did that his answer is the SAME as 1/5 (as is mine)

OpenStudy (unklerhaukus):

yea

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

But i said i need answers not solution

OpenStudy (unklerhaukus):

is -0.2 also valid?

OpenStudy (anonymous):

no it is a positive

OpenStudy (anonymous):

Am closing this now u all should be waiting for another question

OpenStudy (unklerhaukus):

thanks for playing

OpenStudy (anonymous):

No. thank u for participating